# Class-Discriminative Attention Maps for Vision Transformers

### v2

@@ Abstract
Importance estimators are explainability methods that quantify feature importance for deep neural networks (DNN). In vision transformers (ViT), the self-attention mechanism naturally leads to attention maps, which are sometimes used as importance scores for which input features ViT models are focusing on. However, attention maps do not account for signals from downstream tasks. To generate explanations that are sensitive to downstream tasks, we have developed class-discriminative attention maps (CDAM), a gradient-based extension that estimates feature importance with respect to a known class or a latent concept. CDAM scales attention scores by how relevant the corresponding tokens are for the predictions of a classifier head. In addition to targeting the supervised classifier, CDAM can explain an arbitrary concept shared by selected samples by measuring similarity in the latent space of ViT. Additionally, we introduce Smooth CDAM and Integrated CDAM, which average a series of CDAMs with slightly altered tokens. Our quantitative benchmarks include correctness, compactness, and class sensitivity, in comparison to six other importance estimators. Vanilla, Smooth, and Integrated CDAM excel across all three benchmarks. In particular, our results suggest that existing importance estimators may not provide sufficient class-sensitivity. We demonstrate the utility of CDAM in medical images by training and explaining malignancy and biomarker prediction models based on lung Computed Tomography (CT) scans. Overall, CDAM is shown to be highly class-discriminative and semantically relevant, while providing compact explanations.
PyTorch implementation: https://github.com/lenbrocki/CDAM.
Web live demo: https://cdam.informatism.com/.
@@ 1 Introduction
Vision transformers (ViT) are a family of computer vision models based on the transformer architecture [1], that demonstrate comparable or even better performance than convolutional neural networks (CNN) on many object detection and classification tasks [2, 3]. A powerful feature of the transformers is that the attention mechanism provides a built-in explanation of the model’s inner workings, if attention is understood as a weighing of the relevance of individual tokens [4, 5, 6, 7]. While attention maps provide intuitive and inherent interpretability, they fail to be discriminative with respect to a target class. Recognizing the potential of attention maps for explaining a ViT’s inner workings, we have developed class-discriminative attention maps (CDAM), a gradient-based extension that makes them sensitive to class-specific signals based on a classifier or a concept derived from selected samples.
@@
Attention maps are the visualization of the self-attention of the class token query [CLS] in the final attention layer. Attention heads can be visualized separately, and in certain instances, may be related to semantically meaningful concepts (e.g., in language [8]). Recently, ViT trained with self-supervised learning (SSL) methods such as DINO [9] have shown excellent performance leveraging a greater amount of unlabeled data. Attention maps from such self-supervised ViT models provide high-quality semantic segmentations of input images. These attention maps clearly distinguish regions of interest (ROI) from the background with very little noise, with different attention heads often focusing on semantically distinct regions [9]. There have been many improvements to ViT including the introduction of convolution [10, 11] and spatial sparsity [12]. Attention maps, as well as our proposed CDAMs, are applicable to all ViT architectures with self-attention.
@@
The major shortcoming of attention maps, when considering them as an explainable AI (XAI) method, is that they purely operate on the level of the last attention layer and therefore do not take into account any information coming from a downstream task, such as a classifier head. The proposed CDAM overcomes this shortcoming by calculating the gradients of the class score, i.e. the logit value of a target class, with respect to the token activations in the final transformer block (visualized in Fig. 2(a)). Essentially, CDAM scales the attention scores by how relevant the corresponding tokens are for the model’s decision. CDAM therefore inherits the desirable properties of attention maps, but additionally indicates the evidence or counter-evidence for a target class that the ViT relies on for its predictions (examples in Fig. 1). We demonstrate that it is class-discriminative in the sense that the heat maps clearly distinguish the object representing the target class from the background and, importantly, from objects that belong to other classes. Therefore, CDAM is a useful tool for explaining the decision-making process of classifiers with a ViT backbone.
@@
In addition to the classifier-based explanations, we present a concept-based approach that replaces the class score with the similarity measure in the latent space of the ViT, where a concept is defined through selected images. Averaging the latent representations of the example images yields a concept embedding, which is analogous to a concept vector in variational autoencoders (VAE) and related generative models [13, 14]. This allows obtaining a heat map for the target concept the model has never been explicitly trained on, which is of particular interest for self-supervised models. We further demonstrate when computing gradients of a class (Section 3.1) or concept similarity (Section 3.2) score, with respect to the token activations in the final transformer block, smoothing or averaging could be applied to improve performance. Specifically, we introduce two additional variants of CDAM; (1) Smooth CDAM averages over multiple gradients that are computed with additional noise to tokens (inspired by SmoothGrad [15]) and (2) Integrated CDAM takes the intergral of the gradients along the path from the baseline to the token in the final transformer block (inspired by IntGrad [16]).
@@
We conduct several quantitative evaluations focusing on correctness, class sensitivity, and compactness. Correctness is measured by perturbing an increasing amount of input features and measuring the model output [17, 18]. By using the ImageNet samples [19] with multiple objects [20] and applying importance estimators for different classes, we quantify the level of class-discrimination. Sparsity and shrinkage of explanations are measured by the number of near-zero importance scores. Besides vanilla, Smooth, and Integrated CDAM, we also employed attention maps, relevance propagation (RP) [21], partial Layer-wise Relevance Propagation (LRP) [22], Input $\times$ Gradients, SmoothGrad [15], and Integrated Gradient [16] in comparison. Three variants of CDAM outperform other explainability methods on three benchmarks.
@@
Lastly, we have applied CDAM on a ViT fine-tuned on the Lung Image Database Consortium image collection (LIDC) [23]. Two ViT-based models are built for predicting malignancy and biomarkers. CDAM, as well as attention maps, are obtained. CDAM is shown to be highly class-discriminative and semantically relevant, while providing implicit regularization of importance scores.
@@ 2 Related Works
Gradient-based methods operate by backpropagating gradients from the prediction score to the input features to produce feature attribution maps, which is also called saliency maps. Many variants have been proposed, see [24, 16, 25, 15, 26] to just name a few. In particular, CDAM is similar to and inspired by the input×$\times$×gradient method [26], because it involves calculating the gradients with respect to tokens multiplied by their activation (Eqs. 2 and 4). The main difference to most gradient-based methods is that we don’t backpropagate the gradients all the way to the input tokens. CDAM stops backpropagation at the tokens that enter the final attention layer. In that regard, our method is related to GradCam [25] which backpropagates the gradients to the final convolutional feature map in a CNN. GradCam and CDAM therefore share the property that they operate on high-level features which is presumably the reason why they are both sensitive to the targeted class, a feature that many explanation methods do not have [27]. Several methods have attempted to apply gradient-based methods or Layer-Wise Relevance Propagation (LRP) [28] to Transformer architectures [29, 30, 22, 31, 32], mostly in the context of NLP tasks.
@@
Attention has often been used to gain insight into a model’s inner workings [4, 5, 6, 7] since it seems intuitive that this weighing of individual tokens correlates with how important they are for the performed task. There is also an ongoing discussion about whether attention provides meaningful explanations [6, 33, 34]. Furthermore, methods such as “attention rollout” [31] have been proposed that aim to quantify how the attention flows through the model to the input tokens. [21] defines the relevancy matrices, as the Hadamard product of the attention map and the gradient of $f_{c}$ w.r.t. to the attention layer, accumulated at each layer. This relevance propagation (RP) method [21], to the best of our knowledge, shows the highest degree of class-discrimination in ViT models. This is used in our study as one of comparative importance estimators.
@@
Many papers have been published on the problem of evaluating post-hoc explanation methods, leading even to the proposal of meta-evaluations [35]. In a recent overview of quantitative evaluation methods [36] 12 desirable properties have been identified. We primarily focus on three of them, namely correctness, contrastivity, and compactness. In particular, carefully designed occulusion/perturbation studies can help us calculate a fidelity statistics [17], which is also called the symmetric relevance gain (SRG) measure [37]. [38] first demonstrated that many importance estimators, which purport to be sensitive to the internals of a classifier, produce similar importance scores regardless of the relation of inputs and classes in the training data, which they called “the data randomization test”. Generally, if two importance scores with respect to two different classes produce very similar heat maps, the importance estimator is not class sensitive. We quantify this concept in our class sensitivity evaluation benchmark to measure whether an explanation is sensitive to the classes.
@@ 3 Methods
Adaptation of the transformer architecture to computer vision hinges on creating tokens from small patches of the input image [2]. In addition to those image patch tokens, the [CLS] token is added, which is utilized for downstream tasks such as classification. Attention maps are the two-dimensional visualization of the self-attention in the final attention layer with [CLS] as query token, where the attention scores are averaged over all heads.
@@
by how relevant the corresponding tokens $T_{i}$ are for the class or similarity score. $Q$ and $K$ are the query and key matrices, consisting of a set of $d_{k}$ dimensional vectors [1]. We discuss the proportionality of CDAM to the attention scores in Section A.1.
Self-supervised learning (SSL) methods such as DINO can train ViT models whose attention maps are high-quality semantic segmentations of the input, without using labels [9]. Therefore, we utilize a ViT model [Dosovitskiy2020ViT] pre-trained with DINO as a backbone for our experiments and applications. We denote details about the ViT architecture, training schemes, and model performance in the Section 4. Note that the heat maps shown in this article are scaled independently.
@@ 3.1 CDAM for a known class
where $T_{ij}$ is the $j$-th component of the $i$-th token that is fed to the last attention layer (Fig. 2). We found in our experiments that calculating the gradients with respect to the layer-normalized tokens yields far better results than using the unnormalized ones.
This definition is a first-order estimation of the contribution of $T_{i}$ to $f_{c}$ in the sense that it would be exact if $f_{c}$ depended on $T_{ij}$ only linearly. It is inspired by gradient-based explanation methods with the difference that we don’t backpropagate the gradients all the way to the input layer but only to the input tokens to the final attention layer. This allows us to gain insight into the decision-making process of the combination of ViT and classifier based on high-level features. This approach leads to attention maps that are class-discriminative and well-aligned with the semantic relationships of the class and different parts of an image (Fig. 1).
@@ 3.2 CDAM for a latent concept
Another way of probing high-level features learned by the ViT is through the use of a similarity measure $g$ in the latent space of the ViT. We compare the latent representation $l(X)=T_{\texttt{[CLS]}^{\prime}}(X)$ of a sample image $X$ and a concept vector $l_{c}$ [39, 13, 14], i.e. the averaged latent representation of $n$ images that are chosen to represent the concept of interest:
@@
This approach is directly inspired by the concept saliency map [13]. The definition of the importance score in this setting is analogous to Eq. 2
@@ 3.3 Smooth and Integrated CDAM
SmoothGrad (SG) [15] and Integrated Gradients (IG) [16] extended saliency maps [24], which were introduced in the context of CNNs as visualizing gradients of the class with respect to the input features. Essentially, these two modifications sharpen and denoise vanilla saliency maps, with theoretical and practical justifications. Briefly, [15] adds a small Gaussian noise to a sample image $n$ times, resulting in $n$ vanilla saliency maps. SG is an average of those $n$ saliency maps. In IG, instead of adding random noise, [16] uses multiple saliency maps obtained from interpolating between a given sample and a baseline (e.g., a black image).
We have adapted the analogous modifications to CDAM, resulting in two additional variants which we call Smooth and Integrated CDAM. Instead of modifying the input images, we vary the tokens that enter the final transformer block. In Integrated CDAM, we interpolate between a baseline and the tokens themselves, resulting in $n$ maps. In Smooth CDAM, we repeatedly add a small noise to the tokens and again obtain $n$ maps.
@@
where $T_{i}=T_{i}+\mathcal{N}\left(0,\sigma^{2}\right)$ with $\mathcal{N}\left(0,\sigma^{2}\right)$ representing Gaussian noise with standard deviation $\sigma$. For the baseline $\tilde{T}_{i}$ we choose the null vector. We implement these two variants by first propagating an image through the ViT up to the final transformer block, where the activations of the tokens are obtained. Next, the activations are manipulated by either applying the noise or interpolation step and the manipulated tokens are then further propagated to obtain the final ViT output.

### v3

@@ Abstract
Importance estimators are explainability methods that quantify feature importance for deep neural networks (DNN). In vision transformers (ViT), the self-attention mechanism naturally leads to attention maps, which are sometimes interpreted as importance scores that indicate which input features ViT models are focusing on. However, attention maps do not account for signals from downstream tasks. To generate explanations that are sensitive to downstream tasks, we have developed class-discriminative attention maps (CDAM), a gradient-based extension that estimates feature importance with respect to a known class or a latent concept. CDAM scales attention scores by how relevant the corresponding tokens are for the predictions of a classifier head. In addition to targeting the supervised classifier, CDAM can explain an arbitrary concept shared by selected samples by measuring similarity in the latent space of ViT. Additionally, we introduce Smooth CDAM and Integrated CDAM, which average a series of CDAMs with slightly altered tokens. Our quantitative benchmarks include correctness, compactness, and class sensitivity, in comparison to 7 other importance estimators. Vanilla, Smooth, and Integrated CDAM excel across all three benchmarks. In particular, our results suggest that existing importance estimators may not provide sufficient class-sensitivity. We demonstrate the utility of CDAM in medical images by training and explaining malignancy and biomarker prediction models based on lung Computed Tomography (CT) scans. Overall, CDAM is shown to be highly class-discriminative and semantically relevant, while providing compact explanations.
Figure 1: Extending attention maps (AM), the proposed class-discriminative attention maps (CDAM) quantify and visualize input features that are relevant for the target class in Vision Transformers (ViT). We visualize importance scores obtained with CDAM for a linear classifier with ViT-S/8, trained with DINO, as a backbone (Caron et al., 2021). For details, refer to Section 3.1. Orange and blue colors correspond to positive and negative values, respectively.
@@ 1 Introduction
Vision transformers (ViT) are a family of computer vision models based on the transformer architecture (Vaswani et al., 2017), that demonstrate comparable or even better performance than convolutional neural networks (CNN) on many object detection and classification tasks (Dosovitskiy et al., 2020; Wu et al., 2020). A powerful feature of the transformers is that the attention mechanism provides a built-in explanation of the model’s inner workings, if attention is understood as a weighing of the relevance of individual tokens (Mullenbach et al., 2018; Bahdanau et al., 2014; Serrano & Smith, 2019; Thorne et al., 2019). While attention maps provide intuitive and inherent interpretability, they fail to be discriminative with respect to a target class. Recognizing the potential of attention maps for explaining a ViT’s inner workings, we have developed class-discriminative attention maps (CDAM), a gradient-based extension that makes them sensitive to class-specific signals based on a classifier or a latent concept.
@@
Attention maps are the visualization of the self-attention of the class token query [CLS] in the final attention layer. Attention heads can be visualized separately, and in certain instances, may be related to semantically meaningful concepts (e.g., in language (Clark et al., 2019)). Recently, ViT trained with self-supervised learning (SSL) methods (Caron et al., 2021; Touvron et al., 2021; Oquab et al., 2024) have shown excellent performance leveraging a greater amount of unlabeled data. Attention maps from such self-supervised ViT models provide high-quality semantic segmentations of input images. These attention maps clearly distinguish regions of interest (ROI) from the background with very little noise, with different attention heads often focusing on semantically distinct regions (Caron et al., 2021). There have been many improvements to ViT including the introduction of convolution (Wu et al., 2021; Yuan et al., 2021) and spatial sparsity (Li et al., 2021). Attention maps, as well as our proposed CDAMs, are applicable to all ViT architectures with self-attention.
@@
The major shortcoming of attention maps, when considering them as an explainable AI (XAI) method, is that they purely operate on the level of the last attention layer and therefore do not take into account any information coming from a downstream task, such as a classifier head. The proposed CDAM overcomes this shortcoming by calculating the gradients of the class score, i.e. the logit value of a target class, with respect to the token activations in the final transformer block (visualized in Fig. 2(a)). Essentially, CDAM scales the attention scores by how relevant the corresponding tokens are for the model’s decision. CDAM therefore inherits the desirable properties of attention maps, but additionally indicates the evidence or counter-evidence for a target class that the ViT relies on for its predictions (examples in Fig. 1). We demonstrate that it is class-discriminative in the sense that the heat maps clearly distinguish the object representing the target class from the background and, importantly, from objects that belong to other classes. Therefore, CDAM is a useful tool for explaining the decision-making process of classifiers with a ViT backbone. Note that the goal of CDAM is to explain the classifier head and the final transformer block.
@@
In addition to the classifier-based explanations, we present a concept-based approach that replaces the class score with the similarity measure in the latent space of the ViT, where a concept is defined through selected images. Averaging the latent representations of the example images yields a concept embedding, which is analogous to a concept vector in variational autoencoders (VAE) and related generative models (Brocki & Chung, 2019). This allows obtaining a heat map for the target concept the model has never been explicitly trained on, which is of particular interest for self-supervised models. We further demonstrate when computing gradients of a class (Section 3.1) or concept similarity (Section 3.2) score, with respect to the token activations in the final transformer block, smoothing or averaging could be applied to improve performance. Specifically, we introduce two additional variants of CDAM; (1) Smooth CDAM averages over multiple gradients that are computed with additional noise to tokens (inspired by SmoothGrad (Smilkov et al., 2017)) and (2) Integrated CDAM takes the integral of the gradients along the path from the baseline to the token in the final transformer block (inspired by IntGrad (Sundararajan et al., 2017)).
@@
We conduct several quantitative evaluations focusing on correctness, class sensitivity, and compactness. Correctness is measured by perturbing an increasing amount of input features and measuring the model output (Brocki & Chung, 2023c; b). By using the ImageNet samples (Deng et al., 2009) with multiple objects (Beyer et al., 2020) and applying importance estimators for different classes, we quantify the level of class-discrimination. Sparsity and shrinkage of explanations are measured by the number of near-zero importance scores. Besides the proposed vanilla, Smooth, and Integrated CDAM, we also employed 7 different explainability methods in comparison. Three variants of CDAM outperform other methods on quantitative benchmarks.
@@
Lastly, we have applied CDAM on a ViT fine-tuned on the Lung Image Database Consortium image collection (LIDC) (Armato III et al., 2011). Two ViT-based models are built for predicting malignancy and biomarkers. CDAM is shown to be highly class-discriminative and semantically relevant, while providing implicit regularization of importance scores.
@@ 2 Related Works
Gradient-based methods operate by backpropagating gradients from the prediction score to the input features to produce feature attribution maps, which is also called saliency maps. Many variants have been proposed, see (Simonyan et al., 2013; Sundararajan et al., 2017; Selvaraju et al., 2017; Smilkov et al., 2017; Shrikumar et al., 2016) to just name a few. In particular, CDAM is similar to and inspired by the input×$\times$×gradient method (Shrikumar et al., 2016), because it involves calculating the gradients with respect to tokens multiplied by their activation (Eqs. 2 and 4). The main difference to most gradient-based methods is that we don’t backpropagate the gradients all the way to the input tokens. CDAM stops backpropagation at the tokens that enter the final attention layer.
In that regard, our method is related to GradCam (Selvaraju et al., 2017) which backpropagates the gradients to the final convolutional feature map in a CNN. GradCam and CDAM therefore share the property that they operate on high-level features which is presumably the reason why they are both sensitive to the targeted class, a feature that many explanation methods do not have (Rudin, 2019). Several methods have attempted to apply gradient-based methods or Layer-Wise Relevance Propagation (LRP) (Bach et al., 2015) to Transformer architectures (Atanasova et al., 2020; Ali et al., 2022; Voita et al., 2019; Abnar & Zuidema, 2020; Chefer et al., 2021b), mostly in the context of NLP tasks. GradCam for ViT backpropages the gradients to the outputs of the final attention layer (Gildenblat, 2021).
@@
Attention has often been used to gain insight into a model’s inner workings (Mullenbach et al., 2018; Bahdanau et al., 2014; Serrano & Smith, 2019; Thorne et al., 2019) since it seems intuitive that this weighing of individual tokens correlates with how important they are for the performed task. There is also an ongoing discussion about whether attention provides meaningful explanations (Serrano & Smith, 2019; Wiegreffe & Pinter, 2019; Jain & Wallace, 2019). Furthermore, methods such as “attention rollout” (Abnar & Zuidema, 2020) have been proposed that aim to quantify how the attention flows through the model to the input tokens. (Chefer et al., 2021a) defines the relevancy matrices, as the Hadamard product of the attention map and the gradient of $f_{c}$ w.r.t. to the attention layer, accumulated at each layer. This relevance propagation (RP) method (Chefer et al., 2021a), to the best of our knowledge, shows the highest degree of class-discrimination in ViT models. This is used in our study as one of the baseline importance estimators.
@@
Many papers have been published on the problem of evaluating post-hoc explanation methods, leading even to the proposal of meta-evaluations (Hedström et al., 2023). In a recent overview of quantitative evaluation methods (Nauta et al., 2023b) 12 desirable properties have been identified. We primarily focus on three of them, namely correctness, contrastivity, and compactness. In particular, carefully designed occlusion/perturbation studies can help us calculate the fidelity statistics (Brocki & Chung, 2023c), which is also called the symmetric relevance gain (SRG) measure (Bluecher et al., 2024). Adebayo et al. (2018) first demonstrated that many importance estimators, which purport to be sensitive to the internals of a classifier, produce similar importance scores regardless of the relation of inputs and classes in the training data, which they called “the data randomization test”. Generally, if two importance scores with respect to two different classes produce very similar heat maps, the importance estimator is not class-sensitive. We quantify this concept in our class sensitivity evaluation benchmark to measure whether an explanation is sensitive to the classes.
@@ 3 Methods
[…]
Figure 2: Graphical scheme of class-discriminative attention maps (CDAM). (a) CDAMs are obtained by first propagating an image through the transformer blocks and using its latent representation $l=\texttt{[CLS]}^{\prime}$ to infer a class score $f_{c}$ or similarity score $g(l,l_{c})$ (see Eq. 2 or Eq. 4). (b) Detailed view of the final transformer block, adapted from (Dosovitskiy et al., 2020).
@@
by how relevant the corresponding tokens $T_{i}$ are for the class or similarity score. Here $Q_{\texttt{[CLS]}}=T_{\texttt{[CLS]}}W^{Q}$ with $W^{Q}\in\mathbb{R}^{d_{\text{model }}\times d_{k}}$ and $K_{i}=T_{i}W^{K}$ with $W^{K}\in\mathbb{R}^{d_{\text{model }}\times d_{k}}$. In this notation $Q$ and $K$ are the query and key matrices which pack together all the query and key vectors (Vaswani et al., 2017) and the indices, e.g. $K_{i}$, select single query or key vectors.
Self-supervised learning (SSL) methods such as DINO can train ViT models whose attention maps are high-quality semantic segmentations of the input, without using labels (Caron et al., 2021). Therefore, we utilize a ViT model (Dosovitskiy et al., 2020) pre-trained with DINO as a backbone for our main experiments and applications. Furthermore, we demonstrate general applicability of CDAM by using other ViT models, namely pre-trained with Supervised Weakly through Hashtags (SWAG) (Singh et al., 2022), Data-efficient image Transformers (DeIT) (Touvron et al., 2021) and DINOv2 (Oquab et al., 2024). We denote details about the ViT architecture, training schemes, and model performance in the Section 4 and Section A.5. Note that the heat maps shown in this article are scaled independently.
@@ 3.1 CDAM for a known class
where $T_{ij}$ is the $j$-th component of the $i$-th token that is fed to the last attention layer (Fig. 2). We found in our experiments that calculating the gradients with respect to the layer-normalized tokens yields far better results than using the not normalized ones.
This definition is a first-order estimation of the contribution of $T_{i}$ to $f_{c}$ in the sense that it would be exact if $f_{c}$ depended on $T_{ij}$ only linearly. It is inspired by gradient-based explanation methods with the difference that we don’t backpropagate the gradients all the way to the input layer but only to the input tokens of the final attention layer. This allows us to gain insight into the decision-making process of the combination of ViT and classifier based on high-level features. This approach leads to attention maps that are class-discriminative and well-aligned with the semantic relationships of the class and different parts of an image (Fig. 1).
@@ 3.2 CDAM for a latent concept
Another way of probing high-level features learned by the ViT is through the use of a similarity measure $g$ in the latent space of the ViT. We compare the latent representation $l(X)=T_{\texttt{[CLS]}^{\prime}}(X)$ of a sample image $X$ and a concept vector $l_{c}$ (Brocki & Chung, 2019), i.e. the averaged latent representation of $n$ images that are chosen to represent the concept of interest:
@@
This approach is directly inspired by the concept saliency map (Brocki & Chung, 2019). The definition of the importance score in this setting is analogous to Eq. 2
@@ 3.3 Smooth and Integrated CDAM
SmoothGrad (SG) (Smilkov et al., 2017) and Integrated Gradients (IG) (Sundararajan et al., 2017) extended saliency maps (Simonyan et al., 2013), which were introduced in the context of CNNs as visualizing gradients of the class with respect to the input features. Briefly, (Smilkov et al., 2017) adds a small Gaussian noise to a sample image $n$ times, resulting in $n$ vanilla saliency maps. SG is an average of those $n$ saliency maps. In IG, instead of adding random noise, (Sundararajan et al., 2017) uses multiple saliency maps obtained from interpolating between a given sample and a baseline (e.g., a black image).
We have adapted the analogous modifications to CDAM, resulting in two additional variants which we call Smooth and Integrated CDAM. Importantly, instead of modifying the input images, we vary the tokens that enter the final transformer block. In Integrated CDAM, we interpolate between a baseline and the tokens themselves, resulting in $n$ maps. Notice that, since we apply the same method, the axioms of IG still hold with the only difference that they apply to the tokens that enter the final transformer block instead of the input pixels. In Smooth CDAM, we repeatedly add a small Gaussian noise to the tokens and again obtain $n$ maps.
@@
where $T^{\prime}_{i}=T_{i}+\mathcal{N}\left(0,\sigma^{2}\right)$ with $\mathcal{N}\left(0,\sigma^{2}\right)$ representing Gaussian noise with standard deviation $\sigma$. For our experiments, we used $n=50$. For the baseline $\tilde{T}_{i}$ we choose the null vector of 0. We implement these two variants by first propagating an image through the ViT up to the final transformer block, where the activations of the tokens are obtained. Next, the activations are manipulated by either applying the noise or interpolation step and the manipulated tokens are then further propagated to obtain the final ViT output.